OpenStudy (anonymous):

Define a function f : R → R by the formula f(x) = 8x − 3. Prove that f is onto. Let y is in R. Then y + 3 8 is in R , and f y + 3 8 = 8 − 3 = .

OpenStudy (zzr0ck3r):

f(x) = 8x − 3. let y be in R then y = 8((y+3)/8) -3 = f((y+3)/8) so y = f(x) when x = (y+3)/8 qed

OpenStudy (anonymous):

\[Let y \in R. Then \frac{ y+3 }{ 8 } \in R, and f(\frac{ y+3 }{ 8 }) = 8(?)-3 =\]

OpenStudy (zzr0ck3r):

? = (y+3)/8

OpenStudy (zzr0ck3r):

\[f(\frac{y+3}{8})= 8(\frac{y+3}{8})-3= y\]

OpenStudy (amistre64):

\[-\infty\le f(x)\le \infty\] \[-\infty\le (8x-3)\le \infty\] \[-\infty\le 8x\le \infty\] \[-\infty\le x\le \infty\]

OpenStudy (amistre64):

prolly not a "correct" proof tho:)

OpenStudy (zzr0ck3r):

that's cute:)

OpenStudy (anonymous):

>zzr0ck3r, ok, that makes sense now that I see it

OpenStudy (zzr0ck3r):

try and understand what @amistre64 did also because there is good insight there. he showed that if f(x) that is real, we have an x that is real.

OpenStudy (zzr0ck3r):

but any time Im showing onto I let y be in R write the inverse of the function f(x) = y, then composite that with the function

OpenStudy (zzr0ck3r):

compose*

OpenStudy (zzr0ck3r):

got more? These are fun and I need the brush up

OpenStudy (anonymous):

quite a bit actually

OpenStudy (zzr0ck3r):

sweet, ima grab coffee ill brb